# RobotDiffuse: Motion Planning for Redundant Manipulator based on Diffusion Model

### v1

@@ V-E Results and Analysis
Figure 3: Success rate of each method changing over time within ten minutes. Diffusion-based methods can rapidly obtain optimal solutions, whereas RobotDiffuse takes more 6s but gets a better success rate.
Although the success rate of RobotDiffuse may not be significantly higher than that of the baseline methods, the proposed approach stands out for its ability to generate a batch of paths within a limited time frame, regardless of the task or scenario, because of its architectural design. The solution rate reaches as high as 84.9% within 15 seconds, and its success rate surpasses all comparison methods within 5 minutes. After 10 minutes, the success rate is only slightly lower than MPINet, by just 2.5%. Compared to traditional methods, RobotDiffuse shows an 11.6% improvement in success rate.
In terms of path length, our method is 0.819 shorter than the best baseline method based on sampling, demonstrating that motion planning through motion generation can effectively optimize planning results from a global perspective, leading to better planning quality. When compared to MPD, our method is slightly longer by 0.076. Analyzing the output results, MPD produces smoother paths, generally leading to more effective results in most scenarios. However, it fails to make the optimal path in certain situations because its pre-trained large model was developed using a single, unchanging complex scenario. As a result, the model’s output joint motions include unnecessary twists when applied to the simpler scenarios in our dataset.
To summarize, the proposed RobotDiffuse can generate high-quality paths, with shorter path lengths and lower collision rates, which can be more stable in practical applications.
V-F Abation Studies
To validate the effectiveness of our method in terms of point cloud encoders, backbone network of diffusion model, and the introduction of physics guidance, we designed several ablation experiments. All the ablation experiments are trained on ROP datasets and evaluated based on ROP training sets. The results are also shown in Table I.
Effect of PointEnet. In addition to the CAE model, we also tested two other backbones in the step of condition embedding: (1) PointNet model with obstacle point cloud and robot point cloud as inputs; (2) DGCNN model with obstacle point cloud and robot point cloud as inputs. The results indicate that the CAE model is more effective at extracting valuable information from the obstacle point cloud. While PointNet and DGCNN are more mainstream point cloud encoding approaches and perform well on individual tasks [27], they underperform in the tasks presented in this paper. This is likely because these models are supervised and cannot be retrained. In this ablation study, they were jointly trained with the diffusion model, which may have led to information loss during training.
Effect of Diffusion backbones. To validate the effectiveness of incorporating the encoder-only Transformer, we use the traditional U-Net, Transformer, and decoder-only Transformer as predictors in the diffusion model. In the last variant, the condition is introduced through a cross-attention layer. For robotic joint sequence data, the bidirectional attention mechanism in the encoder proves particularly effective, as it enables the model to consider the current joint state concerning both past and future joint states within the sequence. This bidirectional approach allows the encoder to capture long-range dependencies and the global context of joint motion, which is crucial for generating smooth and realistic motion sequences. By integrating information from the entire sequence, the model can more accurately predict the next joint positions or motions. Additionally, relying solely on the encoder further simplifies the overall model structure, reducing computational complexity while enhancing the efficiency and convergence of the modules during training. This streamlined approach contributes to better scalability and performance in real-world applications.
Effect of motion-guide. By comparing the loss function including the $\mathcal{L}_{joint}$, $\mathcal{L}_{joint}+\mathcal{L}_{point}$, and $\mathcal{L}$ without the physical motion guide, this paper demonstrates the effectiveness of each component in the loss function and the guidance. The improvement of RobotDiffuse, based on the predicted sequence itself as proposed in this paper, enhances the success rate in robotic scenarios. Additionally, the incorporation of $\mathcal{L}_{point}$ and $\mathcal{L}_{collision}$ enables the robot to learn attitude error and collision information, which helps the robot generate more effective results. Then, although introducing the diffusion guidance can interfere with RobotDiffuse, leading to an increased collision rate, the overall success rate is higher, primarily because the diffusion guidance addresses challenges in the middle of the motion and allows for better alignment with the target.
[…]
[width=.4autoplay, loop, controls]10fig/movie/movie112
Figure 4: There are two sphere obstacles with a radius of 0.2m in the workspace. The positions of the obstacles are (0, 0.02, 0.63) and (0.57, -0.42, 0.77). Franka Panda moves from the initial pose (1.57, 1.23, 1.68, 1.38, 1.31, 2.85, 1.68) to the target pose (0.21, 1.21, 1.78, 2.45, 1.73, 2.62, 1.52), and ensure the smoothness and safety of the movement, which means that the robot maintains a safe distance of 0.0.5m from the obstacle, and does not overlap with the green area.
V-G Visulaization
For the display of motion planning results, this paper provides a more lightweight visualization scheme, which is based on the Web to set the information of obstacle avoidance tasks and visualize the robot’s motion process, which can provide users with convenient and efficient interaction capabilities and can get started at a low cost, as shown in Fig. 4. This method reads the URDF of the robot based on Three.js and describes the relative motion of the robot’s links based on forward kinematics. For obstacles, it sets the corresponding size of blue geometry to render and uses green geometry to wrap to represent the safe distance.
VI CONCLUSIONS
@@
In this work, we introduced RobotDiffuse, a diffusion-based motion planning approach for redundant manipulators that addresses the challenges of high-dimensional motion planning. By integrating physical constraints through a point cloud encoder and using an encoder-only transformer to capture temporal dependencies, RobotDiffuse enhances both the efficiency and stability of motion planning. Our approach not only overcomes the computational inefficiencies of traditional sampling-based methods but also improves the coherence and quality of the generated trajectories. The proposed method was validated using a complex simulation environment and a large-scale dataset, showing promising results in terms of feasibility and performance. These findings suggest that diffusion models hold significant potential for advancing motion planning tasks, especially in dynamic and complex environments. Future work will focus on further improving the scalability and robustness of the model for real-world applications.

### v2

@@ 5.5 Ablation Studies
Ablation studies confirm the superiority of the CAE model for obstacle feature extraction over PointNet or DGCNN, which likely suffered from information loss during joint training. Regarding the diffusion backbone, the encoder-only Transformer proves most effective, leveraging its bidirectional attention mechanism to capture long-range dependencies and global motion context while streamlining computational complexity for better scalability. Furthermore, the integration of targeted loss components $\mathcal{L}_{\text{point}}$ and $\mathcal{L}_{\text{collision}}$ is pivotal for reducing attitude errors, while the proposed diffusion guidance significantly enhances the overall success rate by ensuring precise trajectory alignment during critical middle stages of motion, effectively balancing a marginal increase in collision frequency with a higher rate of mission completion.
6 Conclusion
@@
In this work, we introduced RobotDiffuse, a diffusion-based motion planning approach for redundant manipulators that addresses the challenges of high-dimensional motion planning. By integrating physical constraints through a point cloud encoder and using an encoder-only transformer to capture temporal dependencies, RobotDiffuse enhances both the efficiency and stability of motion planning. Our approach not only overcomes the computational inefficiencies of traditional sampling-based methods but also improves the coherence and quality of the generated trajectories. The proposed method was validated using a complex simulation environment and a large-scale dataset, showing promising results in terms of feasibility and performance. These findings suggest that diffusion models hold significant potential for advancing motion planning tasks, especially in dynamic and complex environments. In addition, we have created and publicly released a trajectory optimization dataset for non-desktop scenarios. We hope this resource will support and inspire further research in the field. In future work, we plan to explore the application of fast diffusion methods to enhance the efficiency of motion planning. Additionally, we aim to validate our approach on real manipulators with varying numbers of joints to further assess its practicality and robustness.